INAF queue observing, 2020 Jul 07 - Jul 08

Observer: A. Cardwell
Telescope Operator: ??
Astronomers from Italy: Queue
Instrument: LUCI


Due to the covid-19 pandemic situation these observations have been performed by LBTO personel
This is 0.5 night INAF payback

through former agreement with the LBT Italian observer (F.Cusano) via e-mail.

At the sunset the sky is clear.

Night duration (12 degree el twilight) 3:30 - 11:00 UT = 7 h 30m, INAF has 3h 45 min

Observing time: * 3h* * 35min* (96%)

Weather time loss: 0h 0min (0%)

technical time loss * 0 h 10min* (4%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Casasola   LUCI2 H-band, MOS 2/5h 0016-0051  
Telluric   LUCI2 H-band, MOS     for casasola, before and after


Copied from email forwarded by Felice of mer 8 lug 2020 alle ore 11:28

All times are in UT, from the UT night of 20200708.

03:40 Preset to tellruic for IT-2019B-0113, PI Casasola.

03:45 LUCI2 MOS error.

RFBG at 419 tel el at 54

03:48 I have Dave on the phone.

03:58 Issues resolved, it just needed an init of the MOS in the end. Preset sent again (we were on the first step). Acquisition starts with luci2.20200708.0003.

04:05 Offset of -6.23, -8.37 applied.

04:07 The through slit image shows the star to be too close to one side of the slit. Giving it a nudge.

04:11 18 degree twilight.

04:16 Additional manual nudge of +0.3” in x required. Absorbed. Starting science exposures. The RTD was very laggy, but everything else in my x2go session ws fine. I’ll restart the RTD. Science exposures start with 0008. Trace looks good.

04:19 Now on the second slit, trace looks good. Final science exp 0011.

04:20 Preset to main science target. Acquisition exposures start with 0012.

04:24 Acquisition exposures running.

04:31 I had to remove objects 36 and 38 from the calculations, but the errors looked excellent after that.

04:34 Acquisition confirmed in through mask exposure, it looks spot on! Science exposures start with 0016. DX Guider reports 0.8”, DIMM 1”. The sky is clear.

04:51 Subtracting 16 and 17 the traces from the alignment stars are clear, as are the traces from several of the slittlets. DX guider now reports 0.75”, 0.9” from the DIMM. Conditions are perfect for this.

05:12 Starting second pass through the script. 4 repetitions are required in all. DX guider reports 0.7”, 0.9” from the DIMM. Second pass starts with 0025.

05:46 Pass 3 of 4 begins. Conditions remain good, 0.9” from the DX guider, 1.1” from the DIMM the sky remains clear. Wind low, < 10m/s. This pass begins with 0034. 33-32 looks good!

06:21 Fourth and final pass, then another telluric for this program. First exposure of the fourth pass is 0043. 0.9” from the DX guider, 1.1” from the DIMM.

07:00 Finished! Final exp was 0051. Preset to telluric.

07:06 First acq exposure is 0052.

07:10 Initial acq offset is -5.32, -8.04.

07:12 Acquisition looks good. Science begins with 0055.

07:18 Finished with INAF science, handing over to Arcetri a few mins early. Last science exp was 0058.

07:22 Reauth to LUCI1.

07:25 Midpoint of night.
Topic revision: r1 - 22 Sep 2020, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback